OpenStudy (anonymous):

Please solve! \[-4\left| 7-x \right|\le-20\]

OpenStudy (anonymous):

move the -4 to the other side then break it into 2 equations

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

|7 - x| <= 5

OpenStudy (anonymous):

remove the absolute value the same way as in the last problem

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

First divide both sides by negative four. When you divide by a negative number the inequality flips. Therefore, we would end up with\[|7-x| \ge 20/4\] Now breaking up, we get \[(7-x) \ge 20/4 ~~ and ~~ (7-x) \le 20/4\]

OpenStudy (anonymous):

so could it be \[2lex \le12\]

OpenStudy (anonymous):

Another way to write this is as \[-20/4 \le 7-x \le 20/4\]

OpenStudy (anonymous):

oh thank you guys!

OpenStudy (anonymous):

UselessMass, I have to disagree with your answer. The solution I am getting is \[(-\infty, 2] \cup [12,+\infty)\]

OpenStudy (anonymous):

hmm :D

OpenStudy (anonymous):

You flip the inequality anytime you divide or multiply by a negative number.

OpenStudy (anonymous):

kayser, let's review. 1) Divide by -4 and flip the inequality. \[|7-x| \ge 5\] 2) Remove the inequality. Since this is the greater than case, we have to write two separate inequalities. \[7-x \ge 5 ~~ and ~~ 7-x \le -5\]3) Subtract 7 from both inequalities \[-x \ge -2 ~~ and ~~ -x \le -12\]4) Divide through by -1 for both inequalities and remembering the flip the inequalities \[x \le 2 ~~ and ~~ x \ge 12\] In set notation this becomes \[(-\infty, 2] \cup [12, +\infty)\]

OpenStudy (anonymous):

UselessMass, remember that in the case the inequality is greater than the given number, we have to write the inequality as two inequalities. It is impossible to write it as one. http://www.purplemath.com/modules/absineq.htm